Evacute. this is a combination problem.
5C3 x 12C10/11C7
combination method is nCr and the problem above is in nCr form.

Answer by fractalier(6550)

I think you mean "evaluate," yes?
Anyway, 5C3 = (5*4*3) / (3*2*1) = 20
12C10 = 12C2 = (12*11) / (2*1) = 66
11C7 = 11C4 = (11*10*9*8) / (4*3*2*1) = 330
so we have
20 * (66/330) = 4